OpenStudy (anonymous):

simplify (4- squareroot3) ^2

OpenStudy (anonymous):

(4-sqrt3)^2 =16-2sqrt3+3 =19-2sqrt3

OpenStudy (akshay_budhkar):

\[(a-b)^2=a^2-2ab+b^2\] Here \[a=4, b= \sqrt 3\] Just plug in the formula

OpenStudy (anonymous):

can someone answer this one?